My son is having trouble with this and I don't remember how to do this. I will give you a few questions. Simplify the following expressions. Show your work.
3x-7(x+1)
7-(2x+8)
3x+5x-(x+3)
Thanks

Answer by Earlsdon(6294)

Simplify:
1)
3x-7(x+1) Apply the distributive property to the second term.
3x-7x-7 Now combine like-terms (3x-7x = -4x)
-4x-7 ...and that's it!
2)
7-(2x+8) Remove the parentheses and change the signs of the terms inside the parentheses.
7-2x-8 Now combine like-terms (7-8 = -1)
-1-2x and you're done.
3)
3x+5x-(x+3) First remove the parentheses and change the signs of the terms inside.
3x+5x-x-3 Now combine like-terms (3x+5x-x = 8x-x = 7x)
7x-3 Done!